A perpetuity pays $x at the end of each month. The nominal annual rate of interest compounded monthly is i
(12). Calculate the percentage of increase in the value of this perpetuity
if the nominal annual rate of interest compounded monthly decreases by 10%.
1
Expert's answer
2020-10-12T18:56:10-0400

At the rate"i^{12}" the present value of the perpetuity is;

"\frac{x}{i^(12)}"

A the rate "i^{(12)}(0.9)" the present value of the perpetuity is;

"\frac{x}{i^(12)(0.9)}"

percentage increase in the value of perpetuity is found as;


"\frac{\frac{x}{i^(12)(0.9)}-\frac{x}{i^(12)}}{\frac{x}{i^(12)}}=\frac{1}{0.9}-1"

"=11.11111111%"%